OpenStudy (anonymous):

Solve the triangle. Round lengths to the nearest tenth and angle measures to the nearest degree. b = 4, c = 6, A = 85°

OpenStudy (anonymous):

Use the Law of Cosines a^2 = b^2 + c^2 - 2bc * cos(A) Once you have a, you can use the Law of Sines: a/(sin(85)) = b/sinB = c/sinC

OpenStudy (anonymous):

Do you have any other questions?

OpenStudy (anonymous):

thats not any of the answers they have up there sorry .

OpenStudy (anonymous):

This is HOW to do it. Not the answer O.o

OpenStudy (anonymous):

lmao ohhh i was looking for the answer sorry about that

OpenStudy (anonymous):

i kinda figured that was the way to do it

OpenStudy (anonymous):

It's okay ^_^